### v1

# CLIC: Contrastive Learning Framework for Unsupervised Image Complexity Representation

Abstract
As an essential visual attribute, image complexity affects human image comprehension and directly influences the performance of computer vision tasks. However, accurately assessing and quantifying image complexity faces significant challenges. Previous works needed more generalization capabilities and well-labeled datasets to learn image complexity features. However, creating such datasets requires expensive manual labeling costs, and the models inevitably learn about human subjective biases. To address the above problems, we propose CLIC, an unsupervised framework based on contrastive learning, for learning image complexity representations. The method learns image complexity features on unlabeled data, avoiding the high labeling cost. Specifically, we propose a unique positive and negative sample selection strategy to reinforce the differences in complexity features. At the same time, we introduce an image prior-based Complexity-Aware Loss to constrain the learning process of the model. We conducted extensive experiments for verification, and the results show that CLIC can effectively learn the image complexity representation. CLIC obtained competitive results with supervised methods by fine-tuning on IC9600. In addition, CLIC applied to downstream tasks shows significant performance improvements, demonstrating the potential for application in various real-world scenarios. code
1 Introduction
Figure 1: CLIC framework consists of a query encoder and a key encoder with the same structure, and parameter updates are performed by gradient and momentum, respectively. Images in the mini-batch are pre-computed with global entropy as a prior information to construct an auxiliary task. We extract the feature activation energy of the last down-sampling or the last stage with the global entropy of the image to compute the Complexity-Aware loss. In this way, the model is guided to learn the basic features of the image instead of the category or object attributes.
As the complexity of visual tasks increases, accurately assessing and understanding the complexity of images plays a crucial role in downstream task performance (see Appendix A.2). Image Complexity (IC) is a description of the complexity of an image, which is originally a relatively abstract concept that humans can understand [13]. From a subjective point of view, an IC image’s complexity is reflected in its diversity of visual elements, structural complexity, and information density, which together determine the difficulty of human cognition and understanding of the image [24]. Objectively, IC can be regarded as a variety of diverse features in an image, such as color change, texture distribution, and detail representation [44]. Therefore, IC is also one of the crucial attributes to focus on when processing images based on image data-driven computer vision (CV). In recent years, researchers have proposed the concept of computable image complexity to simulate and quantify human visual perception of image complexity through computer algorithms. The quantization and automatic prediction of image complexity have essential roles and applications in several fields, such as image understanding [43], object detection [15], UI design [46], text detection [17], and image enhancement [30].
In order to assess or predict image complexity, some of the earliest work focused on heuristic-based metrics such as information entropy [49], image compression ratio [31], and machine learning methods such as SVMs, random forests, and BP neural networks [9, 19]. However, these algorithms are mainly focus on hand-crafted features, hindering their ability to evaluate the generalization of ICs in practice [12]. Recently, deep convolutional neural networks (CNNs) have demonstrated strong representation and generalization capabilities to explicitly model human subjective perception in a data-driven manner (e.g., image quality evaluation [45], etc.). Several large-scale datasets for evaluating ICs and deep-learning-based IC prediction methods have emerged (e.g., the ComplexityNet [28], ICNet [12], ICCORN [20], etc.). However, building large-scale supervised IC evaluation datasets requires a vast labor cost because of the need to give the ground truth of IC for each image manually. Despite the significant progress made by models such as ICNet, the reliance on manual scoring, with huge labeling costs and subjectivity bias, limits their diffusion in practical applications.
To avoid expensive labeling costs and subjective bias, a very intuitive idea is to utilize unlabeled datasets for IC representation learning, after which the IC score of each image can be computed by fine-tuning with a small number of samples. Self-supervised learning (SSL) [33] generates pseudo-labels by utilizing unlabeled data or creating auxiliary tasks to extract helpful information from the internal structure of the data, which is suitable for situations where labeled data is scarce. SSL can learn valuable features from unlabeled data and reduce the dependence on labeled data. MoCo [23] used momentum encoders and dynamic dictionaries that solve the negative sample insufficiency problem by maximizing the positive sample similarity and minimizing the positive and negative sample similarity through contrastive learning. The success of MoCo triggered extensive subsequent research, such as MoCo v2 [4], SimCLR [3], BYOL [16], and SwAV [1], which pushed forward the development of self-supervised contrastive learning and demonstrated great potential in the field of computer vision.
Self-supervised learning effectively solves the problem of representation learning under unlabeled conditions by constructing pseudo-labels from the data itself. However, extending this approach to the domain of image complexity assessment still faces challenges because of the abstraction and uniqueness of IC features. Driven by the above issues, we propose the contrastive learning framework CLIC (Fig.1) to efficiently learn image complexity features without relying on labeling. Following [23], we construct a query encoder and key encoder, and update the parameters by gradient and momentum. We design a unique positive and negative sample selection strategy for the image complexity representation task. We introduce image prior information to construct a complexity-aware loss, which excludes category interference and guides the model to learn image complexity features accurately. We conducted extensive experiments for validation, and the experimental results show that CLIC can effectively learn the image complexity representation. CLIC fine-tuned in IC9600 yields competitive results with supervised learning. In addition, CLIC applied to downstream tasks shows significant performance improvements, demonstrating the potential for application in various real-world scenarios.
In summary, we make the following contributions:
  • CLIC is proposed as an unsupervised learning framework for image complexity representation, designed to learn the complexity features of images without relying on manual scoring labels.
  • We design a reasonable positive and negative sample selection strategy for the characteristics of image complexity, taking the view closest to the original IC as the positive sample to enhance the contrastive learning for the complexity representation.
  • We propose a complexity-aware loss based on the image, which is a prior information, to optimize the model’s complexity-aware ability by penalizing the difference with the image.
  • Extensive experiments are conducted to demonstrate that CLIC effectively learns image complexity representations. Additionally, CLIC is applied to downstream tasks, resulting in improved performance in both detection and segmentation.
2 Related Works
2.1 Image Complexity Evaluation
Traditional Manual Method. Chipman et al. [7] investigated the complexity factors of visual patterns at a very early stage, and the structural features of similarity transformations and color transformations were able to reduce the perceptual complexity. Ichikawa et al. [26] and Oliva et al. [38] investigated the intrinsic processes underlying human judgments of pattern complexity, exploring how real-world visual complexity is perceived by the human perceptual system. Rosenholtz et al. [41] proposed three measures for the visual clutter problem, and visual clutter is also one of the forms of image complexity. Ramanarayanan et al. [40] used multidimensional scaling MDS to analyze the image complexity. Dimensional scaling MDS to analyze subjects’ judgments of scene complexity. Chikhman et al. [6] used image spatial features, frequency features, and other feature analysis methods to build a computational model to estimate image complexity and compare it with the results given by the observers. Palumbo et al. [39] and Machado et al. [36] both quantified visual complexity based on Image compression algorithms to quantify visual complexity and correlation analysis was used to assess the difference between the quantified results and the subjective complexity.
Intelligent Learning Method. Nagle et al. [37] used a pre-trained convolutional neural network (CNN) to predict complexity ratings for 4000 images, ultimately achieving results that were highly correlated with human observer ratings. Saree et al. [42] explored the extraction of unsupervised information from the intermediate convolutional layers of a CNN, proposing an activation energy metric to quantify visual complexity. Kyle-Davidson et al. [29] used a region neighborhood graph cutting algorithm to capture visually salient regions in an image, computed the local symmetry of an image to extract symmetry features, and built a ComplexityNet to predict image complexity. Feng et al. [12] spent a high cost of labeling to construct an image complexity assessment system that contains 9,600 well-labeled image complexity assessment datasets and provided a baseline model ICNet for predicting image complexity scores and estimating complexity heat maps in a weakly supervised manner.
2.2 Contrastive Learning
InstDisc [48], one of the early approaches to contrast learning, achieves initial representation learning by unsupervisedly treating each image instance as a unique class. MoCo [23] uses momentum encoders to construct a dynamic negative sample pool to solve the problem of insufficient negative samples in mini-batch training. MoCo stores and updates negative samples through a momentum update queue, making the comparison learning perform well in mini-batch training. SimCLR [3] generates the same image with different viewpoints as a pair of positive samples through data augmentation and utilizes large-scale batch processing and projection head networks to enhance the learning effect. Its innovation lies in the data enhancement strategy and large-scale training. SwAV [1] performs comparative learning utilizing clustering, utilizing image clustering results of different perspectives as the comparison target and using online clustering as the comparison target. Results as comparison targets, combining online clustering with comparison learning, enabling the model to perform more efficient representation learning unsupervised. BYOL [16] achieves self-guided learning of representations by alternating the training of two networks (online and target networks). No explicit negative sample pairs are required, demonstrating high-quality representation learning. DINO [50] transfers the knowledge from the teacher’s network to the student’s network using a self-supervised distillation method to achieve efficiency without labeling learning.
3 Method
3.1 Positive and Negative Samples Selection
A judicious selection of positive and negative samples in representation learning can help the model learn meaningful image representations. This meaning is generally defined for a specific task (e.g., similarity of pairs of positive samples for image classification). Therefore, in IC representation learning, we propose a positive and negative samples selection strategy for IC representations to ensure that the model can focus on the complex features of images rather than the content, categories, or object properties. Unlike traditional representation learning, positive and negative sample selection for IC representations needs to be designed around the complexity properties of the images themselves, rather than the similarity of categories or content, or even try to exclude these interfering factors as much as possible.
In the image complexity assessment task, each image needs to be assigned a score to represent its image complexity. Theoretically, the IC scores of any two images will not be equal. Therefore, we consider each image a separate class, i.e., images other than themselves are other classes and are considered negative samples. Some work, such as SimCLR [3], generates different views from the original image as positive sample pairs and images outside the mini-batch as negative samples. We also follow this approach (Fig.2(a)). Nevertheless, the difference is that we do special treatment for the views.
Figure 2: (a) Positive and Negative Samples Selection. We crop the mini-batch image to produce views whose image complexity is close ($\sim$ IC score) to that of the original as positive samples. Outside the mini-batch are negative samples. (b) Small-scale crop and merge. Given an image of size $(h,w)$, the crop size is $(h,w)/c$, $(h,w)/2c$, and $(h,w)/4c$, the corresponding number of crops for each image is $c$, $2c$, and $4c$, respectively. Then, we randomly merge two crops of the same size and their two transformed crops.
Positive sample pairs for the image classification task belong to one class, i.e., the class of randomly crop sub-images from the image is the same class as that of the original image. Moreover, the IC scores of the original image and the resulting view should be as close as possible in IC representation. Therefore, we design crop schemes with different sizes (with data augmentation) and evaluate them with the Pearson correlation coefficient (PCC) [18] between the views and the original image. Finally, the crops with the closest PCC is used as the views and defined as a positive sample pair.
Our framework has no additional definition for negative samples and considers images outside the mini-batch as negative samples while building a dynamic queue to store negative sample features [23]. In addition, we process the original dataset by small-scale crop and merge it in such a way as to introduce local image features for the model learning phase, which will be more beneficial for IC representation learning [12]. Note that small-scale crop and merge (Fig.2(b)) is not just a boost to the original data volume but is more focused on local feature enhancement, allowing the model to learn a localized complexity representation of the image.
3.2 Complexity-Aware Loss with Global Entropy Prior
Contrastive loss, commonly used in representation learning, can distinguish similarities or differences between positive and negative samples. In contrast, similarity computation tends to introduce image category attributes for model learning, which may result in complexity perception errors. Therefore, image complexity representation learning needs to guide the model to focus more on or even prioritize the primary attributes of an image (which should be independent of category, content, and object) when extracting features. We found an essential phenomenon with some simple metrics for image complexity assessment (Tab.1). We randomly sampled 1000 images from the ImageNet training set, computed the IC evaluation results of each simple metric for these samples separately, and used the inference output of ICNet as the ground truth. We found that edge density, UAE [42], and global entropy all have relatively high Pearson correlation coefficients with the ground truth (details in Appendix C), and earlier, some scholars have similarly shown that metrics such as global entropy can evaluate image complexity well [36].
PCC UAE ED CR GE GT
UAE - 0.2912 -0.3594 0.4971 0.5582
ED - - -0.2472 0.3930 0.5034
CR - - - -0.8401 -0.3842
GE - - - - 0.5259
GT - - - - -
Table 1: PCC of image metrics. UAE is unsupervised activation energy. ED is edges density. CR is compress ratio. GT is ground truth.
In addition, Saraee et al. [42] investigated the feature extraction properties of intermediate convolutional layers utilizing deep neural networks capable of capturing both low-level (e.g., edges, texture) and high-level features (e.g., objects and scene content) of an image. They quantize several convolutional feature maps to create a mapping of feature activation energy to image complexity scores in an unsupervised form. Surprisingly, our experiments found the highest PCC between UAE and global entropy, compared to other combinations, and a highly positive correlation with ground truth. A relationship between global entropy and feature activation energy can be established to guide the model in learning a representation close to the ground truth.
Driven by these observations, we consider using an image’s global entropy ge as a prior information to construct an auxiliary task for image complexity representation learning, proposing the image Complexity-Aware Loss (CAL). The goal is to optimize the complexity properties of positive and negative samples. The image complexity characterization task needs to focus the model on image complexity features and use the image prior information to make the model generate representations that minimize the content, category, or object attributes prone to complexity-awareness errors. For each image in the mini-batch, its global entropy is pre-computed using the following form:
$ge=-\sum\limits_{i=0}^{L-1}{p(i)\cdot lo{g_{2}}(p(i))}$ (1)
Where $p(i)$ denotes the probability of gray value $i$, which is obtained by counting the number of pixels in each gray level $L$ (usually 256 for 8-bit gray image) and then normalized, for feature activation energy, we extract the feature maps of each downsampling (every stage in ResNet [21] or SWin [34] series) of the query encoder to calculate the feature activation energy fae and sum it:
$fae=\sum\limits_{s}{(\frac{1}{{h\times w\times d}}\sum\limits_{i,j,k}^{h\times w\times d}{F[i,j,k]})}$ (2)
Where $F$ is the feature map, $h$, $w$, and $d$ are the feature map’s height, width, and depth (number of channels), respectively. $s$ denotes the number of down-sampling. After that, MSE loss is used to construct the image complexity-aware loss $L_{CAL}$ form as follows:
${L_{CAL}}=\frac{1}{N}\sum\nolimits_{i}^{N}{{{(x_{i}^{fae}-x_{i}^{ge})}^{2}}}$ (3)
Where $N$ is the batch size, $i\in[0,N]$. $x_{i}^{fae}$ and $x_{i}^{ge}$ denote feature activation energy fae and global entropy ge, respectively. In addition, following some representation learning methods [23, 4, 3], this paper also considers InfoNCE [47] loss form ${L_{IN}}$ (Eq.(4)) to measure positive and negative sample similarity in terms of dot product while building dictionary look-up problem for contrastive learning.
${L_{IN}}=-\log\frac{{\exp(q\cdot{k^{+}}/\tau)}}{{\exp(q\cdot{k^{+}}/\tau)+\sum\nolimits_{{k^{-}}}{\exp(q\cdot{k^{-}}/\tau)}}}$ (4)
Where $\tau$ is temperature, $q$ is the encoded query. $k^{+}$ is the positive sample key and $k^{-}$ is the negative one. Finally, the loss of CLIC is obtained by combining $L_{CAL}$ as a regular term with InfoNCE:
$L={L_{IN}}+\lambda{L_{CAL}}$ (5)
Where $\lambda$ is a hyper-parameter denoting the weight coefficient of complexity-perceived loss.
3.3 Contrastive Learning Framework for IC Representation
We constructed a framework for image complexity representation based on contrastive learning to learn image complexity representations efficiently. We employ two independent encoders: the query encoder and the key encoder. The query encoder is responsible for extracting complex features from inputs. In contrast, the key encoder processes the inputs and generates a 128 -dimension vector, which is used for contrastive learning with the feature output from the query encoder. Following [23], both encoders share the same network structure, and the query encoder performs gradient updating, and the key encoder uses momentum updating:
${\theta_{k}}\leftarrow m{\theta_{k}}+(1-m){\theta_{q}}$ (6)
Where $\theta_{k}$ is the parameter of the key encoder, $\theta_{q}$ is the parameter of the query encoder, and $m\in[0,1)$ is the momentum factor (default 0.999). The purpose of momentum updating is to help the model maintain a stable perception of image complexity during training by keeping the key encoder’s historical information while avoiding losing important information due to training fluctuations. During training, query and key encoders each receive different transformed views as input. Note that the query encoder receives views with a small amount of data augmentation, while the key encoder’s input has additional data augmentation.
In addition, contrastive learning methods typically require a large amount of data to train the model to obtain the desired representation. Therefore, in this work, we sample from several publicly available datasets [25, 10], utilizing the image global entropy, which has the highest correlation with the ground truth, as the sampling basis. The sampling results show that the global entropy of the images in the training set obeys a uniform distribution ($ge\in(0,1)$). Thus, the true image complexity distribution in the training set is also close to a uniform distribution. The image complexity of the training set is close to the uniform distribution, and the model can learn image complexity representations of different difficulties in a balanced way.
4 Experiments
4.1 Setup
4.1.1 Training Data
Data Collection. Our training images are all derived from a random sampling of Flickr-5B [25] and ImageNe-1M [10], constructing a training set of $\sim$ 1 million images with image entropy obeying a uniform distribution (see Appendix B).
4.1.2 Implementation Details
Training. The query encoder and key encoder are derived from ImageNet pre-training. The batch size is 128 on 4 GPUs. The initial learning rate is 0.03. A total of 200 epochs are trained, with the learning rate reduced by a factor of 10 at the 120th and 160th epochs. The momentum of updating the key encoder is 0.999. Softmax temperature is 0.07. Stochastic Gradient Descent (SGD) is used to optimize the model, where momentum is set to 0.9 and weight decay is set to 0.0001.
Fine-tuning. We fine-tuning the CLIC query encoder on IC9600 with ICNet [12]. The parameters of the CLIC query encoder are frozen during fine-tuning. The batch size is 128 in 4 GPUs. Stochastic Gradient Descent (SGD) is used to optimize the model. The momentum is set to 0.9, the weight decay is set to 0.0001, and the learning rate is fixed to 0.001. Other settings are consistent with ICNet.
4.2 Comparison with Previous Results
We compare CLIC with other methods for evaluating image complexity, both unsupervised and supervised, and we use the same metrics as in [12], i.e., PCC and Spearman’s correlation coefficient (SRCC) [5]. The trained CLIC query encoder is used as the backbone of ICNet, and the parameters are frozen and then fine-tuning on IC9600 train set (Tab.2). Note that ICNet employs a two-branch structure, whereas we only use the CLIC query encoder as a separate branch.
method PCC↑ SRCC↑
unsup. SE 0.534 0.498
NR 0.556 0.541
ED 0.569 0.491
AR 0.571 0.481
UAE 0.651 0.635
super. SAE 0.865 0.860
ComplexityNet† 0.873 0.870
HyperIQA 0.935 0.935
P2P-FM 0.940 0.936
ICNet† 0.947 0.944
ICCORN 0.954 0.951
CLIC Res18 0.890 0.883
Res50 0.893 0.887
Res152 0.899 0.893
Swin-B 0.913 0.908
Table 2: Comparison with previous results. † denotes our reproduction results, and unlabeled results are obtained from the literature [12]. unsup denotes unsupervised learning methods. super denotes supervised learning methods.
The results show that the unsupervised group has the best UAE results, but there is a massive gap between the results and those of the supervised group ICCORN [20]. However, it is worth noting that the unsupervised group methods are traditional hand-designed heuristics. On the other hand, most of the methods in the supervised group are based on deep learning methods and achieve much better performance than the unsupervised methods. This is due to the enormous advantage of deep learning for feature extraction. While CLIC does not perform as well as supervised methods, we aim to address the expensive labeling cost and avoid subjective bias further when we replace the model with a more complex one. Both PCC and SRCC increase to 0.913 and 0.908, which are very close to supervised methods.
4.3 Ablation of CLIC
Regarding the individual members of CLIC, we set up ablation studies to verify their validity. Each case (a $\sim$ e, CLIC, Tab.3) is initialized with an ImageNet pre-trained model, and after training, the CLIC query encoder is fine-tuning on the IC9600 dataset. The results of the table show the validity of the individual members and demonstrate that there is no conflict between them.
case pre-train IC9600
pos scm CAL PCC↑ SRCC↑
(a) 0.759 0.748
(b) 0.796(+.037) 0.781
(c) 0.804(+.045) 0.788
(d) 0.806(+.047) 0.793
(e) 0.839(+.076) 0.826
CLIC 0.893(+.134) 0.887
Table 3: Ablation of CLIC. pre-train denotes pre-training initialization. IC9600 denotes fine-tuning in this dataset. pos denotes the positive sample selection strategy. scm denotes small-scale crop and merge. CAL is complexity-aware loss.
We sampled untrained 1k images from 10 random classes in the ImageNet train set to do t-SNE visualization (Fig.3). ICNet infers these images to get their IC scores. We extracted the output feature of stage 4 in the CLIC query encoder and downscaled them to 2-dimensional space. We labeled their IC scores (divided into ten bands on $(0,1)$) and class id in ImageNet with color, respectively. We find that CLIC successfully clusters samples with closer ICs without clustering similar samples together. This suggests that CLIC can focus on image complexity representation learning without the interference of attributes such as class. We also visualize the results of case (a) (Tab.3) in Appendix D.1.
Figure 3: t-SNE visualization. All feature are taken from the output of stage 4 in CLIC query encoder (ResNet). The complexity score (a) is divided into ten segments. The class id (b) is from 0 to 9.
We select several images from the MS COCO test set and extract the activation map of their convolution in the last layer of the CLIC query encoder (Fig.4). The results present that CLIC is able to disperse hotspots to most regions of the image in cam rather than focusing on individuals. Such results further illustrate that CLIC is able to focus on the primary attributes of an image rather than individual attributes.
Figure 4: Activation map. The case (a) is the same in Tab.3. All feature maps are extracted from stage 4 (ResNet50) of the CLIC query encoder.
4.4 Further Analysis
Figure 5: Positive Views Crop Strategy. The only crop denotes a crop with no data augmentation. The few aug denotes crops with several data augmentation. The more aug denotes crop with more data augmentation.
Positive Views Crop Strategy. In Sec.(3.1), we proposed a special strategy for view crop. We believe that for the image complexity representation task, the complexity between pairs of positive samples should be close to each other, so we carefully design three crop types, randomly sampling 1k images from ImageNet and processing them according to these crop schemes, with the crop size decreasing from 240 to 96. We also use the image global entropy to evaluate these views and calculate the PCC between them and the original image entropy (Fig.5). The results show that all three crop strategy have the most appropriate sizes, and we trained CLIC based on these sizes to get the same results as we expected (Tab.4). First, the augmentation methods become more varied, introducing a rich diversity to the training set, and the performance keeps improving. Second, we made two control sets (case b vs. c, case d vs. e) of suitable and unsuitable crop sizes. We found a difference between them, which is small but validates the effectiveness of our proposed crop strategy.
Contributions of Labels. We compare with supervised ICNet by gradually increasing the number of labels used for learning during CLIC fine-tuning, and the results are presented in Fig.6. Note that ICNet is not fine-tuning but trained from scratch with the same number of labels. First, both methods improve their performance as expected with an increasing number of labels. Then, we observe that CLIC has a PCC close to 0.4 from the beginning with only ten labels, while ICNet is only around 0.1. As the number of labels keeps increasing, CLIC gradually approaches saturation at 5k labels since only the header updates the parameters in fine-tuning. In contrast, the parameters in the feature extraction part are already fixed. However, this trend shows that CLIC, after comparative learning with a small number of labels fine-tuning again, can rapidly increase the image complexity characterization results, proving the effectiveness of our method. In the end, we find that the supervised learning method ICNet is much more accurate than CLIC when a sufficiently large number of labels is reached.
case crop&size PCC↑ SRCC↑
(a) oc 224; oc 224 0.759 0.748
(b) oc 224; fa 144 0.771 0.759
(c) oc 224; fa 224 0.768 0.757
(d) fa 144; ma 144 0.796 0.781
(e) fa 144; ma 144 0.790 0.774
Table 4: Comparison of different positive view selections. The crop&size denotes the crop method and the corresponding size. The oc is only crop, fa is few augmentation crop, and ma is more augmentation crop.
Figure 6: Performance with different labels number.
Complexity-Aware Loss. To verify the role of global entropy in CAL, we chose three cases of compress ratio, edge density, and None for comparison (Tab.5). We find that the performance without CAL is a relatively low 0.759. After setting the prior information in CAL to edge density and global entropy, the PCC improves by 0.22 and 0.47, respectively, which suggests that metrics with higher relevance to the ground truth as a prior are all effective in improving the performance. Similarly, the low (or even negative) correlation between compress ratio and ground truth leads to training failure. The best performance of global entropy is because the hyper-parameter of edge density is not well chosen. Otherwise, both of them can perform close to each other. In addition, we did an ablation of the $\lambda$ of CAL and obtained the optimal value of 0.25 (Tab.6) .
metrics       PCC↑       SRCC↑
None 0.759 0.748
compress ratio       -       -
edge density 0.781(+.022) 0.771(+.023)
global entropy 0.806(+.047) 0.793(+.045)
Table 5: Comparison of different prior metrics. None means no prior metrics.
$\lambda$ 0 0.15 0.25 0.35 0.5
PCC↑ 0.759 0.788 0.806 0.790 0.774
Table 6: Ablation of CAL $\lambda$
4.5 Results of Apply to Downstream Tasks.
Object Detection. We apply our CLIC to Mask R-CNN [22] with several backbones. The train details are the same as MMDetection [2]. The results are shown in Table 6. Overall, the addition of image complexity features improves the detection model’s performance.
Backbone       AP       AP50       AP75
R-50 38.2 58.5 41.4
R-10 40.0 60.5 44.0
X-101-32x4d 41.9 62.5 45.9
R-50† 38.5(+0.3) 59.1(+0.6) 41.8(+0.4)
R-101† 40.2(+0.2) 60.9(+0.4) 44.2(+0.2)
X-101-32x4d† 42.0(+0.1) 62.7(+0.2) 45.8(-0.1)
X-101-32x4d†† 42.3(+0.4) 63.3(+0.8) 46.5(+0.6)
Table 7: Comparison of Mask R-CNN on MS COCO. All models are trained with 1× learning rate schedule. † indicate CLIC is applied. The CLIC query encoder weight of feature map which is add to FPN is 0.5. CLIC is a ResNet18. †† means CLIC is a ResNet152.
Method Mem       FPS       mIoU
FCN 5.6G 19.64 66.53
FCN† 14.5G 15.26(-4.38) 68.51 (+1.98)
FCN†† 14.5G 17.83(-1.81) 68.50 (+1.97)
Table 8: Comparison of FCN on PASCAL VOC 2012. The backbone of FCN is R-50-D8. † indicate CLIC is applied. The CLIC query encoder weight of feature map which is add to FCN head is 0.5. CLIC query encoder is a ResNet101. FCN†† indicates the concurrent processing of CLIC and FCN.
Semantic Segmentation. We apply our CLIC to FCN [35] with R-50-D8. The train details are the same as MMSegmentation [8]. Compared to the FCN, the CLIC-assisted FCN† has a mIoU improvement of 1.98. This demonstrates the auxiliary optimization effect of image complexity on semantic segmentation tasks. We visualize some results of the PASCAL VOC 2012 validation set in Fig.7, where the segmentation results of FCN† show a larger and more accurate coverage of the pixels for each object.
Figure 7: Visualization of FCN with PASCAL VOC 2012 val set.
5 Conclusion
Image complexity assessment can aid in optimizing CV tasks. This paper proposes CLIC, an unsupervised image complexity representation learning framework based on comparative learning. Unlike previous approaches that rely on well-labeled datasets, CLIC can efficiently learn IC features based on unlabeled data. Through extensive experimental validation, the results show that CLIC can learn effective IC representations and performs competitively with supervised learning methods for fine-tuning on the IC9600. In addition, CLIC shows significant improvement in several downstream tasks, demonstrating its potential in real-world scenarios. Our future work is to build more suitable pretext tasks for IC representation learning that are entirely free of interference from attributes such as categories, objects, and so on. Furthermore, we also consider developing IC distillation frameworks as a free lunch to optimize downstream tasks.
Supplementary Material
A Preliminary
A.1 Definition of Image Complexity
Image complexity is more difficult to define. Traditional manual methods can be calculated using image entropy, image compression ratio, etc., but most of them are based on heuristic metrics, which are inaccurate to assess and have insufficient generalization ability. Machine learning methods use supervised learning to model image complexity through high annotation costs, a process that inevitably learns human subjective biases. However, whether it is a manual method or a machine learning method, image complexity modeling can be defined as Eq.(7), where ${x_{i}}$ denotes an image from a dataset, and $f(\cdot)$ denotes the mapping from this image to its image complexity $IC_{i}$, with the range of $IC_{i}$ being 0 to 1.
$I{C_{i}}=f({x_{i}})$ (7)
After analyzing the image complexity of several datasets, we observe that, in general, a dataset’s Image Complexity Distribution (ICD) is close to a normal distribution i.e. $I{C_{i}}\sim N(\mu,{\sigma^{2}})$ , indicating that most of the images are concentrated at a medium level of complexity. As shown in Fig.8, the ICD of the dataset is calculated based on image entropy and machine learning methods, respectively, and the distribution curve belongs to the normal distribution category. It is worth mentioning that most of such datasets with normal characteristic ICDs are natural scene images, such as street, indoor and outdoor, and field. However, some datasets, such as the Cornell Grasp Dataset [27], have high frequencies in their ICD concentrated around 0.17, which is one of the reasons why the learning difficulty of these tasks is low (only when Cornell Grasp Dataset is used for 2D target detection task).
Figure 8: Image Complexity Distribution of Datasets. (a) global entropy of MS COCO [32]. (b) ICD of MS COCO [32]. (c) ICD of PASCAL VOC [11]. (d) ICD of Cornell Grasp [27].
A.2 Impaction of Image Complexity
[…]
IC Range AP AP50 AP75
All 40.5 59.3 43.8
<0.3 41.2(+0.7) 60.3(+1.0) 44.6(+0.8)
>0.7 39.6(-0.9) 58.0(-1.3) 42.6(-1.2)
Table 9: Results of YOLOX-S on MS COCO with different IC range. The training details for YOLOX-S were maintained consistent with the official implementation. All represents the original MS COCO val set, with the IC ranger from 0 to 1.
B Datasets
B.1 ImageNet
ImageNet [10] is a large-scale visual dataset with 1.28 million images in 1000 classes, widely used in computer vision and deep learning. The images’ high complexity and rich visual features, including different lighting conditions, background complexity, angle diversity, etc., make it a high-quality resource for visual representation learning.
B.2 Flikr
Flikr [25] contains $\sim$ 5 billion images of multiple types from Flickr (a popular online photo and video hosting, sharing, and management platform), covering a rich diversity of scenes and content. These images range from common scene types, such as natural scenes from everyday life, portraits, buildings, and city streetscapes, to more creatively expressive content, such as still life and artistic photography. The complexity of the images shows significant differences in the number of visual elements, distribution, texture, light and shadow changes, etc., which is suitable for image complexity analysis and evaluation.
B.3 IC9600
IC9600 [12] is a large-scale benchmark dataset for automated image complexity assessment. The dataset consists of 9600 carefully labeled images, each of which has been labeled in detail by 20 human annotators. The images cover a wide range of domains such as abstraction, advertising, architecture, objects, paintings, people, scenes and transportation.
C Metrics Details
C.1 Unsupervised Activation Energy
Intermediate convolutional layers in deep neural networks act as feature extractors, capable of capturing both low-level (e.g., edges, texture) and high-level features (e.g., objects and scene content) in an image. Driven by this, they define a metric, the activation energy, to quantify the visual complexity of each layer by averaging the activation values (i.e., the feature map) of the convolutional layers.
$UAE=\frac{1}{{h\times w\times d}}\sum\limits_{i,j,k}^{h\times w\times d}{F[i,j,k]}$ (8)
C.2 Edge Density
$ED=\frac{\sum_{i=1}^{H}\sum_{j=1}^{W}\mathbb{I}(\text{Canny}(I(i,j))>0)}{H\times W}$ (9)
Where, $I(i,j)$ is the pixel value at position $(i,j)$ in image $I$. $\mathbb{I}(\cdot)$ is the indicator function, which returns 1 when Canny detects an edge, and 0 otherwise.
C.3 Compress Ratio
$CR=\frac{B}{H(I)}\quad\\ \quad H(I)=-\sum_{x=0}^{255}p(x)\log_{2}p(x)$ (10)
Where, $B$ is the bit depth of the image. $H(I)$ is the entropy of the image $I$, which can be represented by computing the probability distribution $p(x)$ of the image pixel values, $x$ is the pixel value.
D More Ablation
D.1 t-SNE Visualization of Case (a)
We downscale and visualize the stage 4 output features of CLIC query encoder in Fig.3. The results show that CLIC can learn the image complexity representation well without interference from image categories. Here we create a control group and do the same operation for case (a) in Tab.3 and get the result in Fig.9. It is clear that case (a) is not able to learn the image complexity representation effectively compared to CLIC.
Figure 9: t-SNE visualization. All feature are taken from the output of stage 4 in CLIC query encoder (ResNet). The complexity score (a) is divided into ten segments. The class id (b) is from 0 to 9.
D.2 Correlation of Train Data and Complexity-Aware Loss
We used global entropy as a sampling basis to construct the training set and also used it to design the image complexity-aware loss. Intuitively, it is easy to think there is a strong coupling between them. Therefore, we designed multiple sets of controlled experiments to prove that no such strong coupling relationship exists and that other metrics can achieve closer results (Tab.10). We first build a training set to train CLIC using random sampling, and the results obtained for PCC and SRCC are 0.809 and 0.792. Based on this, we apply CAL, and the PCC and SRCC are increased by 0.032 and 0.035, respectively, which shows that the effect of CAL alone is the enhancement of the results is also very obvious. Further, we also use three metrics, namely, compression ratio, edge density, and global entropy, and we collect the data with uniform and Gaussian distributions. The results show that edge density and global entropy are also close. Conversely, the compress ratio has a negative correlation with ground truth, resulting in poorer results. We sampled with Gaussian distribution because more IC scores in IC9600 are distributed between 0.3 and 0.7. However, it may be because most standard Gaussian distributions are concentrated in 0.4 to 0.6. The number of samples of other complexities is too small, which leads to a decrease in the final results (compared to the uniform distribution).
method       PCC↑       SRCC↑
random w/o CAL 0.809 0.792
w/ CAL 0.841(+.032) 0.827(+.035)
uniform compress ratio 0.623 0.604
edge density 0.886 0.881
global entropy 0.893 0.887
gaussian edge density 0.803 0.856
global entropy 0.819 0.812
Table 10: Ablation of train data and CAL. w/ or w/o denotes that complexity aware loss is used or not. The uniform and gaussian denote uniform and gaussian distributions, respectively.
D.3 Small-scale Crop and Merge
Small-scale Crop and Merge ablation results are presented in Tab.11. We find that as $c$ gets progressively larger, the size of the training set rapidly increases from the original 1× to 18×. Further, we find that the PCC and SRCC reach their highest when $c$ is 3, which are 0.906 and 0.900, respectively. However, the PCC and SRCC begin to decay rapidly as $c$ continues to get larger. The reason for this phenomenon is that as $c$ increases the train set is almost overrun by generated images (which can be considered as pseudo images). And too many pseudo images can seriously interfere with the positive value from real data.
$c$ ratio of train set PCC↑ SRCC↑
0 0.890 0.883
2 0.902 0.897
3 11× 0.906 0.900
4 14× 0.881 0.877
5 18× 0.846 0.839
Table 11: Ablation of Small-scale Crop and Merge.
D.4 Image Complexity Assessment
We select images (Fig.10) with different difficulty scores and get their image complexity maps and image complexity scores based on pre-trained CLIC and ICNet fine-tuning. The results show that our framework achieves good results in image complexity evaluation at different difficulty levels.
Figure 10: Visualization of image complexity map and score. We select several images from the IC9600 test set with image complexity scores of 0 $\sim$ 1, and use CLIC query encoder fine-tuning to get their IC map and IC score. The top image of each images pair is the original image, and the bottom image is the IC map. Similarly, the ground truth score is on top and the prediction score is on the bottom.
E Future Work
This paper proposes CLIC, an unsupervised image complexity representation learning framework based on comparative learning. Unlike previous approaches that rely on well-labeled datasets, CLIC can efficiently learn IC features based on unlabeled data. Through extensive experimental validation, the results show that CLIC can learn effective IC representations and performs competitively with supervised learning methods for fine-tuning on the IC9600. In addition, CLIC shows significant improvement in several downstream tasks, demonstrating its potential in real-world scenarios.
Although CLIC can efficiently learn image complexity representations, we believe it is flawed.
  • First, we introduce image prior information as a supervised signal to construct an auxiliary task that helps CLIC learn image complexity representation. However, the contrast loss is still based on InfoNCE, which may inevitably cause CLIC to learn a few categories and object attributes.
  • Second, the trained CLIC needs to parallelize the query encoder with them (backbones) when applying to downstream tasks. This can lead to a loss of efficiency in downstream tasks.
Thus our future work has two parts. We will build a pretext task suitable for image complexity representation learning, which will be more coupled than normal contrast loss. On the other hand, we are going to build a framework for image complexity knowledge distillation, which is expected to migrate the IC knowledge of the query encoder to the downstream task model. This will keep the downstream task efficient while increasing the performance